OpenStudy (anonymous):

5. ABCD is rectangle. Find m

OpenStudy (anonymous):

130 degrees 165 degrees 150 degrees 90 degrees

OpenStudy (anonymous):

;D

OpenStudy (praxer):

it is 130 degree. as in the figure AD II BC and AC is transversal. so, <DAC = < BCA ( corresponding angles) <BCA+<DCA = 90 <DCA=90-<BCA =90-65 =25 Since in a rectangle the diagonals bisect equally so, EC = ED So <EDC=<ECD for the triangle ECD <ECD+<EDC+<DEC=180 <DEC=180-<ECD+<EDC =180-25+25 =130 degree Option A is correct.

OpenStudy (praxer):

@baddiemaddie understood it ???

OpenStudy (anonymous):

Got it, thanks!

OpenStudy (praxer):

welcome.